Transitioning from MPP to SMP:

Experiences with the MIT Ocean-Atmosphere Model

Computation Structures Group Memo 390 February 10, 1997

Chris N. Hill

Andrew Shaw

This report describes research done at the Laboratory for Computer Science and the Department of Earth, Atmospheric, and Planetary Sciences of the Massachusetts Institute of Technology. Funding for the Laboratory for Computer Science is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Of�ce of Naval Research contract N00014-92-J-1310.

1